prove that a gun will shoot three times as high when its angle is 60 as when it is 30 but the bullet will carry the same horizontal distance.
1
Expert's answer
2020-01-10T09:50:02-0500

The horizontal distance is


"R=\frac{v^2}{g}\sin(2\alpha)"

"R_1=\frac{v^2}{g}\sin60=\frac{\sqrt{3}v^2}{2g}"

"R_2=\frac{v^2}{g}\sin120=\frac{\sqrt{3}v^2}{2g}"

Thus,

"R_1=R_2"

"h_1=\frac{v^2}{2g}\sin^230=\frac{v^2}{8g}"

"h_2=\frac{v^2}{2g}\sin^260=\frac{3v^2}{8g}=3h_1"